by TEI TEI

05/24/05 – TE emailed out all the field data collected when helping trying to solve an airflow issue with AHU-3 to BS, AKM and MJ. PDF

05/20/05 – TTC and County discussed flow and pressure issues. TTC stated that trouble could be most likely related to a flex duct being blown off a termination as this is what we have found to cause this type of flow issues in the past.

04/15/05 – TTC stopped at the building on Friday to follow up with overheating of the 3rd floor space. TE tested the airflow into the general spaces and found that almost all areas in questions had too little airflow into the spaces. Some area had as much as 50% less or 100 cfm low airflow. The chilled water temperature and the cold deck air temperature are both right on design numbers – this appears to be an airflow issue at the actual boxes. TTC called A&M and found that they would be willing to give the County HVAC Pro for calibrating the boxes and troubleshooting the system. The serial cable to make the connection cost extra. MJ and TTC talked and MJ will follow up with A&M about getting the software and receiving training in it for this troubleshooting type work. We will start by checking the t-stat operation. The Mitsubishi units will help control this space nicely and the county trades are in the midst of installing one in the CODIS room right now. The electrical permit was approved on 03/21/05 for this work.

03/07/05 – TTC sent an email to A&M asking them to adjust the gain on the boilers for quicker turn down from high fire and to label the DX-9100.

03/03/05 – The County has let ThermalTech balance the pump volume to flows that TE believe will balance the system such that the primary water flow will always be greater than the secondary water flow or in such close proximity that no secondary bypass of water will occur. Overnight, Boiler #1 shutdown on a local controller (either water or high limit) and the plant kept the building warm with the backup boiler all night long. The setpoint varied from its reset value by as much as 15 degrees but the secondary boiler would come on and heat the building up, hence all the floor temperatures were fine this morning. TTC found that Boiler #1 was not running at 8:30 a.m. even though the control system had the boiler commanded on and targeted for 100% firing rate. The problem is that this nuisance trip alarm on the boiler is not reporting as an actual boiler alarm where the horn goes off and this makes it impossible for the DDC system to know that the boiler is in alarm so it must rely on the lag boiler to keep the building warm (this control worked perfectly as demonstrated by the history reports on the DDC system). Changing the flows seems to have worked perfectly as the leaving temperature of the boiler and the hot water secondary temperature have been within 1 or 2 degrees of each other consistently throughout the test. There are still some issues that have to be addressed and I am on it: 1.) The firing rate on the boiler controls seems to overshoot the setpoint to far and to long before backing off the firing rate and this could easily contribute to the boilers shutting down on a local high limit. 2.) The operating limit on the boilers doesn’t seem to shut Boiler #1 down before the High Limit – this forces a manual reset of the boiler as we saw last night. 3.) The actual pump flow for balancing is not know because the new balance points are too far toward the end of the curve and TTC will verify with actual manufacturer what the new flow are – we know the pressure drop at this point. 4.) The pressure gauge ball valves were all left open on the pumps – this causes water to bypass the pumps and decreases he effectiveness of the pumping system. TTC closed these valves and demonstrated that this bypass represented about 10 ft of hd on the secondary pumps and about 3 ft of hd on the primary pumps – these may have been left open when I flow tested the system the first time. 5.) TTC called the boiler manufacturer to discuss alarming points and what and how the low water alarm and the high limit alarm should operate, in conjunction with that, I need to confirm that A&M takes modulation control of the boiler and either the operating limit is in series with the high limit and they still act as safeties or only the high limit is acting as a safety for boiler water temperature. 6.) I believe we discovered that these boilers have a low limit controller that keeps the boiler on low fire when the vessel temperature is below X setpoint (i have to verify with Lynn from Bryan). When the temperature in the system was allowed to go down to 80 degrees last night during testing, we showed that the boiler when started locked into low fire until we took manual control and heated up the system (by the way the local and manual control worked perfectly last night for us and that was a change we had A&M make a couple of month ago when we discovered a power issue between Metasys and DDC). I will work these issues out but at first glance, the changes are working and the results are predictable. I may still have the check valve put into the system once I am convinced the system is operating and working properly.

03/01/05 – TTC has reviewed and calculated various system flows at the Coroner’s Office and presented to Gary Davies. I have targeted and shown reverse flow characteristics as shown by Mike’s and Roger’s field testing. I have found several fixes that can be implemented that require mechanical work in the form of adding a check valve in the bypass location or adding a blind flange in the bypass to prevent reverse flow. Neither of these address the unmatched flow characteristics between the primary and secondary pumps that is necessary to address since the current three way control valve position has all waters diverted into the primary loop under all conditions. TTC will present this possible solution to MJ at the Coroners Office in the next couple of days. With the cold weather we may be able to immediately test the solution under full load conditions.

03/01/05 – MJ contacted TE to discuss boiler rotation appears to automatically rotate. BLR-1, HWP-1 and SHWP-4 are on as of 03/01. We will wait to see what happens on 04/01. A low water alarm does not register on the front end when it happens at this time. We cannot tell from the as-builts how the DDC is connected to the boiler alarm sequence or contact point inside the boiler control panel. Humidifier #3 Burner #3 is locked out and TE contact RCME about this being right or if we were waiting on a couple of parts to fix that. TE stated that as of 09/07/04 we thought the humidifiers were 100% up and running and logged no missing parts features. Steve Garr of RCME confirmed later in the day that indeed everything was 100% working when they left and no additional parts were on order. RCME will assist in discovering why burner #3 is locked out.

02/22/05 – In an ongoing effort to never completely finish a project, we are meeting at the site today to prove that a check valve must be installed in the hot water decoupler between the primary loop and the secondary loop. TTC has already reviewed the P&ID and flows with Gary Davies and there does indeed appear to be a need for a check valve.

11/03/04 – Final Boiler Sequence submitted by A&M. 11×17 dwgs are to follow for O&M Manuals. PDF

10/14/04 – A&M visited the site today and rewired the boiler control panels to operate exactly as TE and Mike Jackson proposed. The Metasys/Local switch will allow the owner to switch between DDC control and local control. Under local control the boiler will either: A) if the boiler panel toggle switch is in “auto”, operate from the local boiler controls (penn controllers) mounted on the boiler shell or B) if the boiler panel toggle switch is in the “hand or local” position, operate directly from the potentiometer mounted on the boiler. The electrical 120V is a result of the combustion air dampers being interlocked to the boiler firing circuit ( TE is proposing to either put a placard indicated where else the power must be shut down or adding a relay that when the boiler main power is off a relay also shuts off this power supply too). A&M remapped the HWS value so that it is correctly displayed on the DDC graphics package.

10/12/04 – A&M has requested to meet Mike Jackson on site to go over his issues with the Boiler plant programming, 120 Volt inside boiler control panel and being able to run the boilers in local control.

09/07/04 – All parties can meet on 09/09/04 at 2:00 pm for startup.

09/01/04 – Geiler onsite reworking humidifier piping. TE to schedule RCME startup.

09/01/04 – Tony has approved state contract. TE working with JCI to provide all documents.

JCI has submitted a revised cost estimate to remove and rework the chiller plant. The new cost is $45,235.00. PDF

Geiler must rework the humidifier installation to separate the intake manifolds.

Automation & Mechanical is supposed to have new boiler programming installed by August 27, 2004.

County must decide to accept/reject JCI new price quote for the chillers.